OpenStudy (anonymous):

[Will give medal] Find the center and radius of the circle. How do I do this? Equation below:

OpenStudy (anonymous):

\[x^2+y^2+8x+7=0\]

OpenStudy (sidsiddhartha):

do u know what is the standard form of a circle?

OpenStudy (anonymous):

Yes. \[(x-h)^2+(y-k)^2=r^2\]

OpenStudy (sidsiddhartha):

then u can try to complete the square first

OpenStudy (anonymous):

4^2=16 \[x^2+8x+16+y^2+\] Not sure what to do next...

OpenStudy (sidsiddhartha):

or what u can do it use anothere orm -- \[x^2+y^2+2fx+2gy+C=0\] this is also a standard form now compare with this equation

OpenStudy (sidsiddhartha):

for this form center=(-f,-g) and \[radius=f^2+g^2-c\]

OpenStudy (anonymous):

let's stick with the first one

OpenStudy (sidsiddhartha):

so u have \[x^2+y^2+8x+7=0 \\and ~the~form ~is--\\x^2+y^2+2fx+2gy+c=0 now ~comparing--\]u'll have \[2f=8 \\f=4 \\and\\ 2g=0\\g=0\] so center is=(-f.-g)=(-4,0) and radius=\[f^2+g^2-c=4^2+0^2-7=16-7=9\]

OpenStudy (anonymous):

you can notice that there is no y term. If we compare our equation with the standard equation.which is this: x^2 + y^2 + 2gx +2fy + c = 0

OpenStudy (anonymous):

yeah what @sidsiddhartha said

OpenStudy (sidsiddhartha):

if u try to complete the square then--\\[x^2+2x*4+4^2+(y-0)^2+7-4^2=0 \\(x+4)^2+(y-0)^2=9\] so center=(-4,0) radius=3

OpenStudy (sidsiddhartha):

and sorry i missed some thing for the first process radius for the previous case=\[\sqrt{f^2+g^2-c}\]

OpenStudy (sidsiddhartha):

got this?

OpenStudy (anonymous):

Yes. Thanks!

OpenStudy (sidsiddhartha):

so u can do any one of them :)